A closed system undergoes a frictionless process from 200 psia and 3 ft3 during which the product pV1.3 remains constant. How much work does the system do if the final pressure is 40 psia?


1
Expert's answer
2021-01-21T03:52:43-0500

Q1

given innitially P1=200psi

v1=3Ft3

=3*1728 inch3

=5189 inch3

and finaly p2=40psi

"\because" PV1.3 = constant

=>P1V11.3=P2V21.3

V21.3=P1V11.3/P2

= 200*(5184)1.3/40

V2=17878.6 inch3


work done = (P2V2-P1V1)/(1-r)

=(40*17878.6-200*5184)/(1-1.3)

1072186.6 lb-in

work done = 121140.82N-M